lolokrat:

Help with Geometry? Find the distance between these given points, using the Pythagorean Theorem. D( 5,-6) E (2,3)

Shadow:

I am sure there is an image to relate the points D and E to us. Can you provide it?

Shadow:

Pythagorean Theorem: \[a^2 + b^2 = c^2\]

lolokrat:

Just a blank graph, I had to plot them myself.

Shadow:

They only give you two points?

lolokrat:

yup

lolokrat:

i typed this up, this is the formula i have on my notes

Shadow:

\[D = \sqrt (x _{2-} x _{1})^2 + (y _{2} - y _{1})^2\]

lolokrat:

yeah thats distance formula i know how to use that one but its telling me to use pythagorean theorem

Shadow:

Pretty sure you would use this one. It's the formula for distance between two points. As you can see, some similarities to Pythagorean Theorem. It is derived from it.

Shadow:

Yeah you are technically using it.